a student mixes two water solutions with an initial temperature of 25.0C to form a final solution with a mass of 65.0g at 30.0C. what is the heat change, in KJ for the reaction?
1
Expert's answer
2012-10-01T10:07:29-0400
Q = c*m*(t2-t1) = 4.18 * 65 g * (30 - 25) = 1.36 kJ